OpenStudy (anonymous):

the difference of two numbers is 3. their sum is 13. find the numbers.. help please :(

OpenStudy (anonymous):

x-y=3 x+y=13 Try to solve...

OpenStudy (anonymous):

i can't D;

OpenStudy (anonymous):

Add the two equations.

OpenStudy (anonymous):

13 and 3?

OpenStudy (anonymous):

x-y + x+y = 3 + 13

OpenStudy (anonymous):

i am sorry i do not get it :/

OpenStudy (anonymous):

x-y is from the first and x+y is from the second 3 and 13 is the same way. Now we get 2x=16 ... x=8 Use the second equation to get y=13-x=13-8=5 The numbers are 5 and 8

OpenStudy (anonymous):

thank you :D